OpenStudy (rizags):

Help Evaluating an Integral:

OpenStudy (rizags):

\[\huge \int\limits_{1}^{\sqrt{3}}\sqrt{4-x^2}-1 dx\]

OpenStudy (thomas5267):

\(x=2\sin(\theta)\)

OpenStudy (rizags):

why?

OpenStudy (anonymous):

Because it helps solving the problem

OpenStudy (rizags):

you are allowed to just make a substitution like that?

OpenStudy (rizags):

oh i see. How would that now affect the upper/lower limits

OpenStudy (thomas5267):

Why not? \[ \int\sqrt{4-x^2}-1\,dx\\ x=2\sin(\theta)\\ dx=2\cos(\theta)\,d\theta\\ \int\left(\sqrt{4-4\sin^2(\theta)}-1\right)2\cos(\theta)\,d\theta \]

OpenStudy (rizags):

So that then becomes \[\int\limits_{}^{}(2\cos(\theta)-1)2\cos(\theta)d \theta\]

OpenStudy (rizags):

now what should i do

OpenStudy (er.mohd.amir):

cos^2 x change into cos 2x and integrate it also change the limits due to substitution

OpenStudy (er.mohd.amir):

cos 2x= 2 cos^2 x - 1

OpenStudy (rizags):

how should i change the limits

OpenStudy (er.mohd.amir):

because in Q variable is x now is theta.

OpenStudy (er.mohd.amir):

x = 2 sin a 1 =2 sin a sin a =1/2 a= 30 degree and when sqr 3 = 2 sin a change it

OpenStudy (rizags):

Is this correct: \[\int\limits_{\pi/6}^{\pi/3}4\cos^2(\theta)-2\cos(\theta)d \theta\]

OpenStudy (er.mohd.amir):

yes

OpenStudy (rizags):

i dont think it is.. i think i went wrong somewhere

OpenStudy (er.mohd.amir):

no its fine. solve it

OpenStudy (rizags):

\[4\int\limits_{\pi/6}^{\pi/3}\cos^2(\theta)d \theta - 2\int\limits_{\pi/6}^{\pi/3}\cos(\theta)d \theta\]

OpenStudy (thomas5267):

\[ \cos(2\theta)=2\cos^2(\theta)-1 \]

OpenStudy (rizags):

that doesnt show up though

OpenStudy (rizags):

i need help with the integral of \[\cos^2(\theta)\]

OpenStudy (rizags):

so far i have \[4\int\limits_{\pi/6}^{\pi/3}\cos^2(\theta) d \theta + (-2\sin(\frac{\pi}{3})-(-2\sin(\frac{\pi}{6}))\]

OpenStudy (thomas5267):

\[ \cos^2(\theta)=\frac{\cos(2\theta)+1}{2} \] Does that make more sense now?

OpenStudy (er.mohd.amir):

if you wants to finf cos^2 t direct use product formula otherwise change to cos 2t

OpenStudy (rizags):

which is \[4\int\limits_{\pi/6}^{\pi/3}\cos^2(\theta) d \theta +(1-\sqrt{3})\]

OpenStudy (rizags):

what is product formula

OpenStudy (er.mohd.amir):

use cos 2 t instant of product formula its make it complected for u.

OpenStudy (rizags):

ok, so once i plug that in then what?

OpenStudy (rizags):

\[4\int\limits_{\pi/6}^{\pi/3}\frac{\cos(2\theta)+1}{2} d \theta\]

OpenStudy (er.mohd.amir):

yes integrate it remember integration of cos ax =1/a (sin ax)

OpenStudy (rizags):

Does that thing become \[2\int\limits_{\pi/6}^{\pi/3}\cos(2\theta)+1 d \theta\]

OpenStudy (rizags):

so: \[2\int\limits_{\pi/6}^{\pi/3}\cos(2\theta) d \theta+2\int\limits_{\pi/6}^{\pi/3}1 d \theta\]

OpenStudy (er.mohd.amir):

yes integrate it

OpenStudy (rizags):

and:\[2\int\limits_{\pi/6}^{\pi/3}\frac{1}{2}d \theta+2\int\limits_{\pi/6}^{\pi/3}\sin(2\theta)d \theta+2\int\limits_{\pi/6}^{\pi/3}1 d \theta\]

OpenStudy (rizags):

wait can i just use this: \[\int\limits_{}^{}\cos^2(x) dx=\frac{1}{2}(x+\sin(x)\cos(x))+C\]

OpenStudy (rizags):

ok so its \[\frac{\pi}{3}+(1-\sqrt{3})=\frac{\pi+3-3\sqrt{3}}{3}\]

OpenStudy (rizags):

woah nice thanks!

OpenStudy (thomas5267):

\[2\int\limits_{\pi/6}^{\pi/3}\cos(2\theta)\,d \theta+2\int\limits_{\pi/6}^{\pi/3}1 \,d \theta=[\sin(2\theta)+2\theta]_{\frac{\pi}{6}}^\frac{\pi}{3}\]

OpenStudy (rizags):

so thats where it comes from.. Ok

OpenStudy (thomas5267):

Did you integrated something twice?

OpenStudy (rizags):

yea i made a mistake and plugged the integral back into the operator